SOUL COM Night #4 - UT 02-Oct-2020

OSA: D.G. Huerta

ISp: G. Bechetti

Mtn Manager: B. Bingham

ISA: B. Rothberg (Secret bunker)

AO: J. Power, D. Mller (Tucson)


Night Plan:

  1. OSCO from start to finish
  2. 2nd OSCO target
  3. Pal 10 - Bin 2 off-axis, J, Ks
  4. M52 - Bin 2 on-axis, K, H, and J (if time)
  5. GRB host galaxy (Bin 4 on-axis)
  6. Performance Testing (Gopt, Modal, Bin4)

Problems:

  • Bin 4 work - Moon may be affecting performance, ran into problems with camera limits being hit
  • Still show an offset between catalog and WFSmag values

Night Log:

UT 00:15 - LUCI GUis started, 2 test darks taken

L1 0001-0002

UT 00:20 - D.G. Huerta starting up WFS, setting shell

UT 01:15 - All-sky a bit hazy (think Stephen King), but roof camera doesn't show same thing - opening enclosure

OSCO

UT 01:30 - pointing and collimation checks complete
  • UT 01:40 - collecting OSCO HD 229143
  • AO Info: Catalog Rmag 7.52, OCAM mag 8.8, bin 1, freq 1700, 58 average pupil intensity, 45.8 photons/subapp EL 72
  • -D.Miller testing telemetry ollection simultaneous to OSCO
  • UT 01:46 - moving nto HD 216399
  • AO Info: Catalog Rmag 7.54, OCAM mag 9.2, bin 1, freq 1700, 60 average pupil intensity, 33.8 photons/subapp EL 41
  • UT 01:57 - Third OSCO target BD+38 3564
  • UT 01:58 - sending preset
  • AO info: Catalog Rmag 7.51, OCAM mag 8.8, bin 1, freq 1700, 40 average pupil intensity, 48.2 photons/subapp EL 84

Pal 10

UT 02:13 - sending preset

-run AO in auto for preset. tests if offsets cause problems

-inserted pause after first line in case need to shift Fieldstop

L1 0063 -aligning N30 fieldstop

L1 0064-0065 - single DIT to verify fieldstop - looks good

AO Info: Catalog Rmag 12.06, OCAM mag 13.4, bin 2, freq 914, 37 average pupil intensity, 5.4 photons/subapp EL 74

AO Ref star: objectName="1085-0406795" ra="289.52158" dec="18.58337"

Gains pre Optical gain 12.54, 0.41, 1.05, gains post optical gain 2.10, 0.07 0.17 with Optical gain 0.18, NCPA trigger disabled

UT 02:25 - continuing script

UT 02:36 - DIMM 0.91"

UT 02:36 - PSF J-band offaxis 0.05" (top left closest to AO Ref star). 0.06" bottom right furtherst from AO ref star, ,PSFs are elongated along the direction of the AO Ref star

L1 0080- loops opened? IIF says we are closed - looks like the AO Ref star walked off the OCAM

L1 0081 - stars look like hot dogs

pause inserted after L1 0081

UT 02:42 - AO GUIs frozen or slow

-D. Miller using PSFGUI secondary GUI to walk the star back

-IIF thinks the AO loops are still closed so GCS is not sending corrections

-Slopes are not being calculated - (star is back) - this looks to be that the BCU has hung, need to restart BCU

UT 02:47 - D. Miller opened the loop, slopes appeared. Now will close the loop.

FIX: open the loop (slopes start calculating), then walk the star back onto the camera, and then close the loop

LUCI Rotator is < 0. Instrument re-rotator turned off in AO system - corrections are poor

This will be an IssueTrak

UT 02:52 - back on track, continuing script

UT 02:59 - Ks (60sec) shows 0.058" top left to 0.10" bottom right

IT 03:07 - 0.9 DIMM

L1 0066-0079 OK

L1 0082-0090 OK

PLEASE NOTE 0086-0090 Are not usable - LUCI rotator < 0 so re-rotator in AO is off - corrections poor

luci1.20201002.0084.fits 1.4094493333
luci1.20201002.0085.fits 0.6528245
luci1.20201002.0086.fits -0.2794354722
luci1.20201002.0087.fits -1.1649451944
luci1.20201002.0088.fits -2.0328796389
luci1.20201002.0089.fits -2.8789730556

SEDOND FIX - PCS calculate rotator for amount of time on preset, if its going negative do an Unwrap

M52 - K,H,J (if time)

UT 03:14 - sending preset

Guider before loops closing ~0.6"

UT 03:19 - DIMM 0.79"

AO Info: EL 51 LUCI Rotator 128, Catalog Rmag 12.35, OCAM mag 13.7, bin 2, freq 839, 27 average pupil intensity, 5.6 photons/subapp
Gains pre Optical gain 10.77, 0.71, 0.88, gains post optical gain 1.28 0.08 0.10 with Optical gain 0.124, NCPA trigger disabled

UT 03:24 - AO offset timeout after 40 seconds

-D. Miller went to bayside stage (X) moved it 0.1 - after 2nd attempt, on 3rd attempt the nudge helped it move

Off-field guiding - FWHM 0.81"/0.73"

L1 0095- safe skipped on first few frames

- the cause is that when we went to the off field active guiding picks up and pulls the star 2" offset on the AO cameras. Solution is to avoid putting offs where active optics / guding would function

L1 0090- L1 0097 - NO NCPA

L1 0098 - L1 0115 NCPA ON

UT 04:00 - DIMM 0.70"

UT 04:12 completed
GRB 200826A

UT 04:12 - sending preset to PSF star

safe skipping on optimize gain, opened loop and then closed it again (auto mode)

- change magnitude to 16.51 in script from catalog value

Bin 4 starts at 16.5, set to 16.6 to ensure Bin 4

will put in intervention mode

UT 04:21 - sending preset

Guider/WFS before closing loop show 0.85", DIMM shows 0.94" EL60

-might be railed on camer lens

UT 04:28 - sending oringal PA=0 script but with AORef mag forced to 16.6 to force it into Bin 4

Guider/WFS before handover to AO shows 0.8"/0.6"

UT 04:32 - DIMM 0.96"

-moving Xpositoin of AO camera to see if it moves back...otherwise looks like camera is railed in X again

UT 04:38 - new script PA=180 manually set Rmag=16.51

-failed to find guidestar

UT 04:42 - doing a pointing check

UT 04:44 - sending preset again

UT 04:46 - DIMM says 0.93"

LUCI Rotator 271 EL65

UT 04:55 - still trying to close the loop -

UT 04:56 - no NCPA applied - closing the loop

loop closed but LUCI is stuck, no timeout - skipped to next line

-Telesvc shows abort wait for preset can be hit meaning LUCI doesn't think the AO preset didn't complete

UT 05:00 - Start AO pressed, failed. Clicked Abort wait for Preset in Telsvc and it continued in LUCI GUI, offset went through then everything hung

-CCD 47 no longer shows the star

-D.Miller trying to move the star back onto CCD47, opening loop then closing again

AO Info: provided Rmag 16.61 (Actual catalog 15.27), OCAM mag 16.3, bin 4, freq 491, 21 average pupil intensity, 4.1 photons/subapp
Gains pre Optical gain 2.7, 1.0, 0.1,GOPT would not apply. Optical gain of 1.000, Aux loop process down and will not come back up

UT 05:11 - AO loop now back, sending LUCI offset

L1 0116-0121 - only doing 6 images on PSF star (enough S/N)

UT 05:21 - moving to GRB field

Guider/WFS before loop show 0.5"

DIMM shows 0.9"

UT 05;24 - resending preset

LUCI Rotator 95 degrees

AO info: provided Rmag 16.61 (Actual catalog 15.06), OCAM mag 16.1, bin 4, freq 491, 26 average pupil intensity, 4.4 photons/subapp
Gains pre Optical gain 1.9, 1.0, 0.1,GOPT would not apply. Optical gain of 1.000, Aux loop process down and will not come back up

AO Ref star FWHM 0.1" --> summed 90sec test image does not show anything

UT 06:00 - seeing from guider (offset in patrol field ) 0.61"

UT 06:10 - DIMM 0.94"

UT 06:30 - DIMM 1.05"

Skipping a few offs (2 or 3)

L0151 off active guiding Guider/WFS show 0.72"/0.89" EL 84 LUCI Rotator=111 degrees

WFS and AdSec control GUI out of sync

Opening the loops fixed the GUIs, now in sync

open loop, stop and restart WFS arbitrator

then resend the preset

L1 0122-0151

L 0152 - BAD - safe skipping - AO Ref star walked off OCAM?Will send a new script with only a few offs

UT 06:50 - EL 87

Transits at 88.8 EL

L1 0153-IQ not great, safe skipped on next frame aborted exposure

Sending new version WITH the corrected RMag to match what it was before

Guider/WFS 0.5"

We'll close the loop again and go for another 15 min (technially night center is 00:10, we've been handing over at 00:00 last three nights)

LUCI Rotator = 244 degrees, but AO camera position is at limit (Y)

UT 07:14 - safe skipping

-why so many camera limits being hit in Bin4? multiple rotator angles 0,90,270, 245?

UT 07:15 - resending preset - Arcetri wants to check on why we keep hitting these camera limits in Bin 4 - this is a technical check for them

-they will need to do work on daytime for Bin4 (TBD that wasn't completed before run)

-x position on camera maxed again

LUCI Rotator = 256

UT 07:21 - center pupil timed out
Handing over to Arcetri

-will do a very quick 30min test of NS as proof of concept for Arcetri - no PSF stars, no collecting data for science, purely technical

-Fortuna was not planned for tonight (too low after midnight) - made a short test script for Eugenia using the standard method - no psf star - don't have time to find one at the moment

AO42 00:40:56.23 +31:22:56.5 RADEC 2000.0 J2000 -39.0 -335.0 12.80 R 2.00 B_V 0.85

Problem with connections to AO computers on the mountain
AO Info: Catalog Rmag 12.8, OCAM mag 13.1, bin 2, freq 978, 47 average pupil intensity, 8.1 photons/subapp Spectral Type M4
Gains pre Optical gain Gains before optical gain 9.99, 0.52, 1.02, with optical gain applied 1.50, 0.08, 0.15
UT 08:02 collectin data
UT 08:24 - taking a sky +40" DETX absolute
Eugenia

UT 08:29 - sending preset

Rmag sent R=12.0, OCAM reporrts 13.4

Gain should be higher for extended objects

DIMM 0.92"

Not sub-framing, max counts is 21k in 2.51 sec

Doing instead 2.51sec x 26 NDIT in a cube. Dither is up and down, should have fixed it for a proper dither, but didn't expect it to be faint.

Guessed Rmag 12 (Vmag 12.189), OCAM mag 13.4, bin 2, freq 904, 38 average pupil intensity, 6.5 photons/subapp

L1 0174 - L1 0194

DIMM 0.93"

UT 09:22 - completed Eugenia script
Star 11 R=15.7

UT 09:29 - sending preset

LUCI Rotator 185 degrees

UT 09;32 - missed GS, doing a pointing check

UT 09:40 - acquiring in Bin 3 but moving to Bin 4

AO Info: El 61, Rot 188, Catalog Mag 15.3, OCAM16.6, bin 4, freq 435, average pupil intensity 29, 3.1 photons/subapp

- lost light in OCAM. vent doors were closed - OCAM dropped to 17.1

-moonlight helped fix the missing flux, prediction was actually off.

UT 10:04 - loops opened

UT 10:05 - DIMM shows 1.24"

EL 61 LUCI Rotator 201

J. Power recovering to move star back into pupils

-started safe skipping

UT 10:14 - resending the preset

DIMM 1.32"

UT 10:24 - loops closed, taking images. Peak counts about half what they were before.

AO Info: Resent preset for star 11, vent doors closed this time: Catalog mag: 15.3, OCAM 16.7, DIMM 1.24", el 60, rot 209, bin 4 freq 457, 18 average pupil intensity, 3.0 photons/subapp

UT 10:32 - looks to have been a seeing bubblle, DIMM is down to 1.1"

UT 10:40 - offsetting for a background iamge

0during 2nd skyframe loops open and AO Ref star wandered off

-issue may be the BCU2K

Star 12 R=15.7

UT 10:54 - moving to Star 12 R=15.7

UT 10:57 DIMM = 1.1"

AO Info: El 58, rotator 222, Catalog mag 15.3, OCAM mag 16.9, bin 4 freq 340, average pupil intensity 16, 1.6 photons/subapp. Gains pre GOPT 3.00, 1.00, 0.10, Gains after GOPT 0.34, 0.11, 0.01.

Seeing 1.03 on the DIMM
UT 11:39 - having some issues with Gopt, processes seem stuck


Star 13 R=16.3

UT 11:49 - sent but wrong flat got loaded

UT 11:53 - Resending preset

guider/wfs show 0.64"/0.6" before loops close

UT 11:58 - DIMM 1.02" EL 52

UT 12:04 - DIMM showing 1.39"

AO info:

UT 12:07 - still waiting for LUCI exposure to be taken, AO loops are closed

-this is the 2nd time that LUCI hold in limbo after AO loops are closed

UT 12:10 - -LUCI timeout just kcked in - it was several minutes between RunAO completing and LUCI timing out

AO Info: el 53, rot 242,, Catalog mag 16.4, OCAM mag 17.3, bin 4, freq 266, average pupil intensity 15, 3.0 photons/subapp. Optimize gain overestimated ott and ho1 again. and luci1 out of sync after done and did not recognize AO preset done

UT 12:25 - DIMM 0.75"@zenith

AO 238 R=12.59

UT 12:26 going to AO 248 06:03:54.09 +26:08:55.8 RADEC 2000.0 J2000 343.0 -543.0 12.59 R 2.78 B_V 0.85
using FeII filters LUCI Rotator 384 degrees

Guider/WFS 0.49"/0.59"

-closing loop failing - E. Pinna trying to diagnose

AO Info: AO248 el 83 rot 384,, Catalog mag 12.59, OCAM mag 13.1, bin 2, freq 924, sky is getting pretty bright. The realtime pupils disappeared when starting close loop. No processes obviously down,

Guider WFS has no spots - way too bright.
UT 12:40 - calling it a night
UT 12:42 - LUCI-1 in safe config - LUCI-2 was not used.

-- BarryRothberg - 02 Oct 2020
Topic revision: r2 - 02 Oct 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback